A major cab company in Chicago has computed its mean fare from O'Hare Airport to the Drake Hotel to be $3.35, with a standard deviation of $26.70 Based on this information, complete the following statements about the distribution of the company's fares from O'Hare Airport to the Drake Hotel.
A. According to Cheb. Therom what percentage of fares fall between $20.00 and $33.40.
B. According to Cheb Therom 84% of all fares fall between $XX.XX and $XX.XX
C. Suppose the disatribution is bell shaped, according to the emperical rule, 99.7% of the fares fall between $xx.xx and $xx.xx
D. Suppose dist. is a bell curve, according to the emperical rule, what percentage of fares fall between $20.00 and $33.40

Answer by stanbon(75887)

A major cab company in Chicago has computed its mean fare from O'Hare Airport to the Drake Hotel to be $26.70, with a standard deviation of $3.35
-------
Based on this information, complete the following statements about the distribution of the company's fares from O'Hare Airport to the Drake Hotel.

A. According to Cheb. Thereme what percentage of fares fall between $20.00 and $33.40.
From $20 is 2 std below the mean; 33.40 is 2 std above the mean.
Cheb says you have at least (1-(1/2)^2) = 0.75 = 75% of data between
those limits.
----------------------
B. According to Cheb Therem 84% of all fares fall between $XX.XX and $XX.XX
If (1-(1/k^2) = 0.84
1/k^2 = 0.16
1/k = 0.4
k = 10/4 = 5/4 standard deviations.
(5/4)(3.35) = 2.51
----
Lower limit: $26.70-2.51 = $24.19
Upper limit: $26.70+2.51 = $29.21
----------------------------------

C. Suppose the distribution is bell shaped, according to the emperical rule, 99.7% of the fares fall between $xx.xx and $xx.xx
Find the z-value with a left tail of 0.3/2= 0.15% = 0.0015
invNorm(0.0015) = -2.9677
Lower Limit: -2.9677*3.35+26.70 = $16.76
Upper Limit: +2.9677*3.35+26.70 = $36.64
========================================
Cheers,
Stan H.

D. Suppose dist. is a bell curve, according to the emperical rule, what percentage of fares fall between $20.00 and $33.40